OpenStudy (anonymous):

The equation 2x^2-x+2=0 has roots a and B. Find the value of a^3+B^3.

OpenStudy (zepp):

Assuming a and B are real solutions Then a+B = -b/a = --1/2 = 1/2 and a*B = c/a = 2/2 = 1

OpenStudy (anonymous):

we know that, the sum of roots of a quadratic eq. is given by - A+B = -b/a so, here, A+B = 1/2 And, Products of roots = c/a = 2/2 = 1 Now ,\[A^{3}+B^{3} = (A+B)^{3}-3AB(A+B)\] putting the values, \[A^{3}+B^{3} = (1/2)^{3}-3(1)(1/2)\] \[A^{3}+B^{3} = (1/8-3/2)\] \[A^{3}+B^{3} = -11/8\]

OpenStudy (zepp):

Meh, everyone, stop typing faster than me :(